4) Fiwrt Corporation manufactures and sells stainless steel coffee mugs. Expected mug sales 4 Fiwrt (in units) for the next three months are as follows: 

October November December 36,000 Budgeted unit sales 30,000 10p00 34,000 


Firt likes to maintain a finished goods inventory equal to 30% of the next month's estimated sales. How many mugs should Fiwrt plan on producing during the month of November? 

A) 34,300 mugs B) 36,000 mugs C) 26,800 mugs D) 35,400 mugs
Answer #1

Mugs to be produced in the month of November = 30% of December's Expected Sales in Units + 70 % of November's Expected Sales in Units

= 30% * 34,000 + 70% * 36,000

= 10,200 + 25,200

= 35,400 units

Hence the correct answer is D) 35,400 mugs
Answer #2

month of November 

=  sales = 36000 unit 

+ Closing stock = 10200 (34000*30%)

- opening stock = (10800) (36000*30%)

= 35400 mugs 

Answer = D    


answered by: Pooja kasera